OpenStudy (anonymous):

Help !!! Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x - 4 y = 3x + 7 y = 3x - 7 y =1/3 x + 2 y =1/3 x - 2

OpenStudy (anonymous):

Use the point slop form and then convert it to y-intercep form

OpenStudy (anonymous):

so y=mx+b . but i dont know which letters go with the numbers

OpenStudy (anonymous):

Point slope form = \( \huge y – y_1 = m(x – x_1)\) Put the point (3, 2) into that equation. \( \huge y – y_1 = m(x – x_1)\) \( \huge y – 2 = m(x – 3)\) Since we know it is going to be parallel, we know that it is going to have the same slope of y = 3x - 4. The slope is 3 so we have \( \huge y – 2 = 3(x – 3)\) Now transfor this \( \huge y – 2 = 3(x – 3)\) into y-intercept form y = mx + b Can you finish ?

OpenStudy (anonymous):

y-2= 3x+-3?? Im sorry im so confused

OpenStudy (anonymous):

You are on the right track. \( \huge y – 2 = 3(x – 3)\) Now distribute the 3 in the () \( \huge y – 2 = 3(x – 3)\) \( \huge y -2 = 3x - 9\) Now add 2 to each side of the equation. \( \huge y -2 + 2 = 3x -9 + 2 \) Can you finish?.

OpenStudy (anonymous):

y=3x-7

OpenStudy (anonymous):

Correct!!!

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

So B

OpenStudy (anonymous):

Yep!!!

OpenStudy (anonymous):

Thank you soo much for helping me and being so patient

OpenStudy (anonymous):

Anytime!!! That is what I am here for :-D

OpenStudy (anonymous):

Thank you for learning and working hard!!!

OpenStudy (anonymous):

Only thanks to you